OpenStudy (anonymous):

Multiply [x-(6-4i)][x-(6+4i)]

OpenStudy (anonymous):

FOIL it out so you get x^2 -(6 - 4i)x -(6 + 4i)x + (36 + 16)

OpenStudy (anonymous):

much easier than it looks

OpenStudy (anonymous):

simplify from there

OpenStudy (anonymous):

first of \((a-bi)(a+bi)=a^2+b^2\) so your constant is \(6^2+4^2=36+16=52\)

OpenStudy (anonymous):

then \[-6-6=-12\] (the i's will cancel so ignore them) and you get \[x^2-12x+52\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x^2-12 x+52

OpenStudy (anonymous):

thats the answer @sate?

OpenStudy (anonymous):

thanks a lot